A small city park consists of a rectangular lawn surrounded on all sides by a 330m(squared) border of flowers 2.5m wide. Find the area of the lawn if entire park is 5 m longer than it is wide.
Answer by mananth(16946)

let width of total park be x m
length = x+5 m
..
Area = x(x+5)= x^2+5x
...
uniform width is 2.5 m
lawn width = x-5 m
lawn length = x+5-5
..
area = x(x-5)= x^2-5x
..
x^2+5x-x^2+5x =330
10x = 330 m
/10 x = 33 width of lawn
length = 33-5 = 28 m
Area of lawn = 28*33= 924 m^2
...
m.ananth@hotmail.ca